OpenStudy (yanasidlinskiy):

For the function \(\normalsize{f(x) = (8-2x)^2}\) ,find \(\normalsize{f-1}\) . Determine whether \(\normalsize{f-1}\) is a function.

OpenStudy (astrophysics):

Well to find the inverse, here are some steps: To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (yanasidlinskiy):

Ok. Let me see what I can get. If not, I'll need serious help. Cuz this is what one of my answer choice looks like: \[f ^{-1}(x) = \frac{ 8\pm \sqrt{x} }{ 2 }\]

OpenStudy (yanasidlinskiy):

I seem not to understand this problem.

jimthompson5910 (jim_thompson5910):

This page helps you determine if a function has an inverse http://www.uiowa.edu/~examserv/mathmatters/tutorial_quiz/geometry/vertandhorizlinetests.html I've highlighted the portion in question (see attached)

OpenStudy (astrophysics):

Yes, that seems to be correct, to find out if it's a function you can do the horizontal line test.

OpenStudy (yanasidlinskiy):

So, pretty much all I have to do is like find out if it intersects or not?

OpenStudy (yanasidlinskiy):

What about solving though?

OpenStudy (zzr0ck3r):

you dont need to do that.

OpenStudy (zzr0ck3r):

if its a function then no one x can point to more than one y

OpenStudy (zzr0ck3r):

how many answers are there for f(2)?

OpenStudy (zzr0ck3r):

hint: the \(\pm\) gives it away.

OpenStudy (zarkon):

you mean \(f^{-1}(2)\)

OpenStudy (zzr0ck3r):

yes \(\uparrow\)

OpenStudy (zzr0ck3r):

I thought that would make it easier, maybe not.:(

OpenStudy (zarkon):

that's how I would do it. you have the inverse...use it

OpenStudy (zzr0ck3r):

@YanaSidlinskiy do you understand? Also, I think you are making it hard on yourself with the normal size stuff `\(\normalsize{f}\)` -> \(\normalsize{f}\) `\(f\)` -> \(f\)

OpenStudy (zzr0ck3r):

unless you change to something different, and need to go back, there is no point.

OpenStudy (yanasidlinskiy):

Sorry..I was afk. Wow! It's crazy let me think.

OpenStudy (yanasidlinskiy):

Ok. I have 4 multiple choice answers. It's half and half. There's 2 that "is a function" and "is not a function"

OpenStudy (zzr0ck3r):

so if f^-1 is a function, then you are not allowed to get two answers to f^-1(x) for any x. what is f^-1(2)?

OpenStudy (yanasidlinskiy):

The other one is: \[\pm \sqrt{\frac{ 8+x }{ 2 }}\]

OpenStudy (zzr0ck3r):

we dont need the options, we know the answer.

OpenStudy (zzr0ck3r):

better yet, how many answers does f^-1(2) give?

OpenStudy (yanasidlinskiy):

Umm..Like 1 or 2. Right?

OpenStudy (astrophysics):

\[f ^{-1}(x) = 4 \pm \frac{ \sqrt{x} }{ 2 }~~~ f ^{-1}(2) =?\]

OpenStudy (yanasidlinskiy):

Ok, I was being stupid on this question. Anyways I got the answer:D Thanx for all your help!!!:D Everyone!!!!!!!!!!!! I seriously do appreciate it!

OpenStudy (yanasidlinskiy):

\(\huge\cal\color{Lime}{Thank~you!!!!!!!!}\)

OpenStudy (zzr0ck3r):

is it a function?